OpenStudy (anonymous):

4x^2+5x-6 I am having trouble solving this. If someone can show me the method to solve, it would be greatly appreciated!

OpenStudy (anonymous):

Quadratic Formula

OpenStudy (anonymous):

Find the roots using the quadratic formula then insert the roots into (x - root)(x - root) to factorize

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

4x^2+5x-6=0 x^2+(5/4)x-3/2=0 (x+5/8)^2=3/2+25/64=121/64 (x+5/8)=11/8 or -11/8 x=3/4 or -2

OpenStudy (across):

If you want to factor that, then you need to find a combination of numbers such that (ax+b)(cx+d)=4x^2+5x-6. In this case, (4x-3)(x+2)=4x^2+5x-6.

OpenStudy (anonymous):

\[4x^2+5x-6 = 0 \implies 8x + 5 = \pm \sqrt{121} \implies 8x = \pm11 -5 \implies x = -2 or 3/4\]

OpenStudy (anonymous):

x = -2,3/4

OpenStudy (anonymous):

\[(2+x) (-3+4 x) \] This means that either \(x=-2 \text{ or } x = \frac34 \)

OpenStudy (anonymous):

Thanks Everyone!!!

OpenStudy (across):

Haha, he's like: "Okay, okay, yes, I got it, guys! Could you leave now?" :P You're welcome.

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@Across Lol